OpenStudy (anonymous):

The first term of an arithmetic sequence is represented by 3x+2y. The third term is represented by 7x. Write the expression that represents the second term.

OpenStudy (anonymous):

Please show me the steps D:

OpenStudy (anonymous):

an=a1+(n-1)d then a3=a1+2d 7x=3x+2y+2d 4x-2y=2d d=2x-y now a2=a1+d=3x+2y+2x-y=5x+y

OpenStudy (anonymous):

the answer is 5x + y though :(

OpenStudy (anonymous):

ohhhh!!!

OpenStudy (anonymous):

never mind, thank you!!